OpenStudy (anonymous):

Greg combined different weights of fertilizer, soil, and compost to make 29 kilograms (kg) of potting soil for some plants. The weight of fertilizer was x kg. The weight of soil was 2 kg more than three times the weight of fertilizer, and the weight of compost was half the weight of fertilizer. This is represented in the equation below. x + (3x + 2) + x = 29 What was the difference in the weights of soil and compost Greg combined? Answer 17 kg 3 kg 6 kg 14 kg

jimthompson5910 (jim_thompson5910):

"The weight of fertilizer was x kg" so x = weight of fertilizer ------------------------------------------------------- "The weight of soil was 2 kg more than three times the weight of fertilizer" So 3x+2 = weight of soil ------------------------------------------------------- "the weight of compost was half the weight of fertilizer.", So x/2 = weight of compost

jimthompson5910 (jim_thompson5910):

To find our answer, we have to solve for x first x + (3x + 2) + x/2 = 29 2x + 2(3x + 2) + x = 2*29 2x + 6x + 4 + x = 58 9x + 4 = 58 9x = 58-4 9x = 54 x = 54/9 x = 6 So the weight of the fertilizer is 6 kg

jimthompson5910 (jim_thompson5910):

Now plug x = 6 into 3x+2 and simplify 3x+2 3(6)+2 18+2 20 So the weight of the soil is 20 kg

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

And finally, plug x = 6 into x/2 to get x/2 6/2 3 so we know that the weight of the compost is 3 kg

jimthompson5910 (jim_thompson5910):

The last part is to subtract the weights of the compost from the soil to get: 20 - 3 = 17 which is the final answer you're welcome